OpenStudy (anonymous):

Verify the identity (1/1+sinx)+(1/1+cscx)=1

OpenStudy (anonymous):

On the Left Hand Side, add the fractions. Do this by common denomitorizing them. So multiply the left by (1 + csc x)/(1 + csc x) and the right one by (1 + sin x)/(1 + sin x). Then add the fractions. The denominator should be (1 + sin x)(1 + cos x). The numerator will be 1 + csc x + 1 + sin x. So show that the numerator and denominator actually are equal.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ok what if I have ((1-2sin^2(θ))/(sin(θ)cos(θ)))=cot(θ)-tan(θ) I have (1-2sin^2θ)/(sinθ(cosθ/sinθ)=(cotθ-(sinθ/cosθ)? where do I go from there?